OpenStudy (anonymous):

PLEASE HELP QUICKLY WILL GIVE A MEDAL Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are 36 and 2304, respectively

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Choices an = 9 ⋅ 4n+1 an = 9 ⋅ 4n+3 an = 9 ⋅ 4n-1 an = 9 ⋅ 4n-2

Directrix (directrix):

Courtesey of @pinterestaddict15 Since T(n)=T(1)r^(n-1) => T(1)r=36--------(1) T(1)r^4=2304--------(2) (2)/(1)=> r^3=2304/36=> r=4 From (1), get T(1)*4=36=> T(1)=9 => the general term T(n)=9*4^(n-1)

Directrix (directrix):

@faariat You are welcome.